# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ 4 Crossing densities of graphons
We will first introduce the rectilinear crossing density of a graphon, but before that, we need some preliminaries. For any four (not necessarily distinct) points $x,y,z,w\in\mathbb{R}^{2}$, let $I(x,y,z,w)=1$ if $\text{conv}\{x,y\}\cap\text{conv}\{z,w\}\neq\emptyset$ and $I(x,y,z,w)=0$ otherwise. We claim that $I$, as a function from $(\mathbb{R}^{2})^{4}$ to $\mathbb{R}$, is measurable. Indeed, let
$X=\{(x,y,z,w)\in(\mathbb{R}^{2})^{4}:\ |\{x,y,z,w\}|\leq 3\ \text{or}\ x,y,z,w\ \text{are contained in a single line}\}.$
The restriction of $I$ to $X$ can easily be seen to be measurable. Hence, it suffices to show that $I^{-1}(1)\backslash X\subset(\mathbb{R}^{2})^{4}$ is measurable. For every positive integer $m$, consider a grid-like subdivision of the plane into interior disjoint squares of side length $1/m$, which we denote by $S_{m,1},S_{m,2},\dots$ in any order. Let $A_{m}\subset\mathbb{N}^{4}$ be the set of all $4$-tuples $(a,b,c,d)$ such that $I(x,y,z,w)=1$ whenever $x\in S_{m,a},\ y\in S_{m,b},\ z\in S_{m,c},\ w\in S_{m,d}$. Write
@@
By the discussion above, this integral is well defined. The function $f$ can be thought of as a straight-line drawing of $W$; in fact, the definition of $\overline{\mathop{\mathrm{cd}}}(W,f)$ is essentially the continuous (and normalized) analog of the definition of $\overline{\mathop{\mathrm{cr}}}(G,\mathcal{D})$ given in Section 2. We divide by $8$ to make up for the fact the each crossing is counted eight times by the integral. The rectilinear crossing density of $W$ is defined as
@@
The definition of the crossing density of a graphon is somewhat more intricate. Let $\mathcal{C}$ denote the family of all curves $c:[0,1]\rightarrow\mathbb{R}^{2}$ that are either simple or constant [^8] Up to this point, we had used the term curve to refer to the image of a continuous function from an interval to the plane. From here on, both the function and its image will be referred to as curves. A simple graphon drawing is a function $\mathcal{D}:[0,1]^{2}\rightarrow\mathcal{C}$ such that the following properties hold:
@@
    For any $x,y\in[0,1]$, $\mathcal{D}(x,y)(0)=\mathcal{D}(x)$.
@@
    For any $x_{1},x_{2},y_{1},y_{2}\in[0,1]$, if the images of $\mathcal{D}(x_{1},y_{1})$ and $\mathcal{D}(x_{2},y_{2})$ are not equal (as subsets of $\mathbb{R}^{2}$), then they have at most one point in common.
For a graphon drawing $\mathcal{D}$ and $x,y,z,w\in[0,1]$, let $I_{\mathcal{D}}(x,y,z,w)=1$ if the curves $\mathcal{D}(x,y)$ and $\mathcal{D}(z,w)$ have at least one point in common, and $I_{\mathcal{D}}(x,y,z,w)=0$ otherwise. Let $\mathcal{D}_{\mathcal{W}}$ denote the family of all simple graphon drawings $\mathcal{D}$ such that $I_{\mathcal{D}}$ is measurable. Given a graphon $W$ and a graphon drawing $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$, write
@@
Remark: We feel the need to clarify that, even though we make a slight abuse of notation by using each of $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ to denote both a graph parameter and a functional on $\mathcal{W}_{0}$, it is not necessarily true that $\mathop{\mathrm{cd}}(G)=\mathop{\mathrm{cd}}(W_{G})$ or $\overline{\mathop{\mathrm{cd}}}(G)=\overline{\mathop{\mathrm{cd}}}(W_{G})$ (see Theorem 4.3 below). Also, it is not completely evident from the definitions that $\overline{\mathop{\mathrm{cd}}}(W)\leq\mathop{\mathrm{cd}}(W)$, but this inequality does hold for every $W\in\mathcal{W}_{0}$.
At this point, it is natural to ask whether for every graphon $W$ there is an $f\in\mathcal{F}$ such that $\overline{\mathop{\mathrm{cr}}}(W)=\overline{\mathop{\mathrm{cr}}}(W,f)$, or a $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$ such that $\mathop{\mathrm{cd}}(W)=\mathop{\mathrm{cd}}(W,\mathcal{D})$ (in other words, can the infimum be substituted by a minimum in the above definitions?). Unfortunately, we do not know the answer to either of these question. The following result, however, is an immediate consequence of the Banach-Alaoglu theorem.
@@
Hence, setting $\overline{\mathcal{I}}$ to be the set formed by all elements of $A$ and all sequential limit points of $A$ with respect to the weak* topology does the job. In the same way, we can show that the set $\mathcal{I}$ which consists of all elements of $B$ and all of its sequential limit points has the required properties. ∎
For every $W\in\mathcal{W}_{0}$, let $\overline{I}_{W}\in\overline{\mathcal{I}}$ and $I_{W}\in\mathcal{I}$ denote two functions that attain the minimums in the statement of Lemma 4.1. The questions above can now be restated as follows. Given a graphon $W$, can the functions $\overline{I}_{W}$ and $I_{W}$ be chosen so that $\overline{I}_{W}=I_{f}$ for some $f\in\mathcal{F}$ and $I_{W}=I_{\mathcal{D}}$ for some $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$?
One of the main goals of this section is to show that the crossing and rectilinear crossing densities are continuous with respect to $d_{\square}$ (see Theorem 4.4). In order to accomplish this, we first require a weaker version of this result.
Theorem 4.2.
@@
$\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W),\ \lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W).$
@@
The tensor product of any two graphons $W_{1}$ and $W_{2}$ is the function $W_{1}\otimes W_{2}:[0,1]^{4}\rightarrow[0,1]$ defined by $(W_{1}\otimes W_{2})(x,y,z,w)=W_{1}(x,y)W_{2}(z,w)$.
Consider a sequence of graphons $W_{1},W_{2},\dots$ that converges to $W\in\mathcal{W}_{0}$ almost everywhere. Clearly, the sequence $W_{1}\otimes W_{1},W_{2}\otimes W_{2},\dots$ converges to $W\otimes W$ almost everywhere. Choose a subsequence $W_{1}^{\prime},W_{2}^{\prime},\dots$ of $W_{1},W_{2},\dots$ with $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n}^{\prime})=\liminf_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})$ and, for every $n$, let $f_{n}$ be an element of $\mathcal{F}$ such that $\overline{\mathop{\mathrm{cd}}}(W_{n},f_{n})\leq\overline{\mathop{\mathrm{cd}}}(W)+1/2^{n}$. By Banach-Alaoglu, we can pass to a subsequence $W_{s(1)}^{\prime},W_{s(2)}^{\prime},\dots$ of $W_{1}^{\prime},W_{2}^{\prime},\dots$ such that $\overline{I}_{f_{s(1)}},\overline{I}_{f_{s(2)}},\dots$ converges to some $I\in\overline{\mathcal{I}}$ in the weak* topology. This way,
@@
for every positive integer $n$, and so the expression on the middle converges to
@@
as $n$ goes to infinity. On the other hand, now a standard analysis argument shows that the sequence $(W_{s(1)}\otimes W_{s(1)})\cdot\overline{I}_{f_{s(1)}},(W_{s(1)}\otimes W_{s(1)})\cdot\overline{I}_{f_{s(1)}},\dots$ converges to $(W\otimes W)\cdot I$ in the weak* topology. Putting everything together, we get that
@@
$\overline{\mathop{\mathrm{cd}}}(W_{n})\leq\overline{\mathop{\mathrm{cd}}}(W_{n},\overline{I}_{W})=\frac{1}{8}\int_{[0,1]^{4}}(W_{n}\otimes W_{n})\overline{I}_{W}(x_{1},x_{2},x_{3},x_{4})\prod_{i=1}^{4}dx_{i}$
for every $n$. Since the expression on the right converges to $\overline{\mathop{\mathrm{cd}}}(W,I_{W})=\overline{\mathop{\mathrm{cd}}}(W)$ as $n$ goes to infinity, we arrive at $\limsup_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\leq\overline{\mathop{\mathrm{cd}}}(W)$, but we also had $\liminf_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\geq\overline{\mathop{\mathrm{cd}}}(W)$; this can only occur if $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W)$. The proof of $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$ is analogous. ∎
Next, we establish a formal connection between the crossing densities of graphons and the crossing densities of graphs.
@@
For every edge weighted graph $G$ on $n$ vertices, we have that
[…]
@@
Let $v_{1},v_{2},\dots,v_{n}$ denote the vertices of $G$ and $I_{1},I_{2},\dots,I_{n}$ be the corresponding subintervals of $[0,1]$ (see the definition of $W_{G}$ in subsection 2.2). The argument is similar to the one in the proof of Theorem 3.4. We begin by showing the inequalities about rectilinear crossing densities.
Given any $f\in\mathcal{F}$, we get a drawing $\mathcal{D}_{f}$ of $G$ as follows: For every $v_{i}$, choose uniformly at random an element $x_{i}$ from $I_{i}$; the point $f(x_{i})$ will represent $v_{i}$. Draw every edge of $G$ as a segment joining its endpoints and, if necessary, perturb the vertices slightly so that no three of them lie on a line and no three edges have a common interior point. Note that the vertex perturbation step can be carried out without increasing the crossing number. A straightforward computation shows that
$\mathbf{E}[\overline{\mathop{\mathrm{cd}}}(G,\mathcal{D}_{f})]\leq\overline{\mathop{\mathrm{cd}}}(W_{G},f).$
Since this holds for every $f\in\mathcal{F}$, we get that $\overline{\mathop{\mathrm{cd}}}(G)\leq\overline{\mathop{\mathrm{cd}}}(W_{G})$.
Suppose now that we are given a rectilinear drawing $\mathcal{D}$ of $G$. This drawing induces an $f_{\mathcal{D}}\in\mathcal{F}$, which is defined by $f_{\mathcal{D}}(x)=p_{i}$, where $i$ is such that $x\in I_{i}$ and $p_{i}$ is the point that represents $v_{i}$ in $\mathcal{D}$. The value of $\overline{\mathop{\mathrm{cd}}}(f_{\mathcal{D}},W_{G})$ is upper bounded by $\overline{\mathop{\mathrm{cd}}}(G)+1/n$, where the $1/n$ accounts for the the $4$-tuples of points in $[0,1]$ which contain two elements of some interval $S_{i}$ (these $4$-tuples are similar to the bad $4$-tuples in the proof of Theorem 3.4). This yields $\overline{\mathop{\mathrm{cd}}}(W_{G})\leq\overline{\mathop{\mathrm{cd}}}(G)+1/n$.
@@
$\mathbf{E}[\mathop{\mathrm{cd}}(G,\mathcal{D}^{\prime})]\leq\mathop{\mathrm{cd}}(W_{G},\mathcal{D}),$
@@
Finally, for every drawing $\mathcal{D}$ of $G$ we obtain an element $\mathcal{D}^{\prime}$ of $\mathcal{D}_{\mathcal{W}}$ in the following way: For every $x,y\in[0,1]$ with $x\in I_{i},y\in I_{j}$, set $\mathcal{D}^{\prime}(x,y)$ to be the curve that represents the edge $(v_{i},v_{j})$ whenever $i\neq j$, and set it to be a constant otherwise (this constant is precisely the point that represents $v_{i}$ in $\mathcal{D}$). As with the rectilinear crossing densities, this graphon drawing satisfies $\mathop{\mathrm{cd}}(W_{G},\mathcal{D}^{\prime})\leq\mathop{\mathrm{cd}}(G)+1/n$, and the result follows. ∎
Theorems 3.5 and 4.3 can now be combined to obtain the following key result.
@@
The crossing density and the rectilinear crossing density are continuous with respect to the cut norm $d_{\square}$. Furthermore, if $G_{1},G_{2},\dots$ is a convergent sequence of edge weighted graphs and the number of vertices of $G_{n}$ goes to infinity with $n$, then $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(G_{i})$ exists and is equal to $\mathop{\mathrm{cd}}(W)$. The same is true with $\overline{\mathop{\mathrm{cd}}}$ in place of $\mathop{\mathrm{cd}}$.

### v4

@@ 4 Crossing densities of graphons
4.1 Defining the crossing and rectilinear crossing densities of graphons
Our first goal in this section will be to introduce the rectilinear crossing density of a graphon. Let us start with some preliminary definitions:
For any four (not necessarily distinct) points $x,y,z,w\in\mathbb{R}^{2}$, let $I(x,y,z,w)=1$ if $\text{conv}\{x,y\}\cap\text{conv}\{z,w\}\neq\emptyset$ and $I(x,y,z,w)=0$ otherwise. We claim that $I$, as a function from $(\mathbb{R}^{2})^{4}$ to $\mathbb{R}$, is measurable. This follows from the fact that $I(-1)$ can be thought of as an algebraic subset of $\mathbb{R}^{8}$. Alternatively, we can let
$X=\{(x,y,z,w)\in(\mathbb{R}^{2})^{4}:\ |\{x,y,z,w\}|\leq 3\ \text{or}\ x,y,z,w\ \text{are contained in a line}\}.$
The set $X$ is measurable, and the restriction of $I$ to $X$ can easily be seen to be measurable too. Hence, it suffices to show that $I^{-1}(1)\backslash X\subset(\mathbb{R}^{2})^{4}$ is measurable. For every positive integer $m$, consider a grid-like subdivision of the plane into interior disjoint squares of side length $1/m$, which we denote by $S_{m,1},S_{m,2},\dots$ in any order. Let $A_{m}\subset\mathbb{N}^{4}$ be the set of all $4$-tuples of distinct indices $(a,b,c,d)$ such that $I(x,y,z,w)$ must equal $1$ whenever $x\in S_{m,a},\ y\in S_{m,b},\ z\in S_{m,c},\ w\in S_{m,d}$. Write
@@ 4.1 Defining the crossing and rectilinear crossing densities of graphons
By the discussion above, this integral is well defined. The function $f$ can be thought of as a straight-line drawing of $W$; in fact, the definition of $\overline{\mathop{\mathrm{cd}}}(W,f)$ is essentially the continuous (and normalized) analog of the definition of $\overline{\mathop{\mathrm{cr}}}(G,\mathcal{D})$ given in Section 2. Note that we divide by $8$ to make up for the fact the each crossing is counted eight times by the integral. The rectilinear crossing density of $W$ is now defined as
@@
The definition of the crossing density of a graphon is somewhat more intricate. Let $\mathcal{C}$ denote the family of all curves $c:[0,1]\rightarrow\mathbb{R}^{2}$ that are either simple or constant [^8] Up to this point, we had used the term curve to refer to the image of a continuous function from an interval to the plane. From here on out, both the function and its image will be referred to as curves. A simple graphon drawing is a function $\mathcal{D}:[0,1]^{2}\rightarrow\mathcal{C}$ such that the following properties hold:
@@
    For any $x,y\in[0,1]$, $\mathcal{D}(x,y)(0)=\mathcal{D}(x)$ and $\mathcal{D}(x,y)(1)=\mathcal{D}(y)$.
@@
    For any $x_{1},x_{2},y_{1},y_{2}\in[0,1]$, if the images of $\mathcal{D}(x_{1},y_{1})$ and $\mathcal{D}(x_{2},y_{2})$ have nonempty intersection (as subsets of $\mathbb{R}^{2}$), then their intersection is connected.
For a simple graphon drawing $\mathcal{D}$ and $x,y,z,w\in[0,1]$, let $I_{\mathcal{D}}(x,y,z,w)=1$ if the curves $\mathcal{D}(x,y)$ and $\mathcal{D}(z,w)$ have at least one point in common, and $I_{\mathcal{D}}(x,y,z,w)=0$ otherwise. Let $\mathcal{D}_{\mathcal{W}}$ denote the family of all simple graphon drawings $\mathcal{D}$ such that $I_{\mathcal{D}}$ is measurable. Given a graphon $W$ and a drawing $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$, write
@@
Remark. We feel the need to clarify that, even though we make a slight abuse of notation by using each of $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ to denote both a graph parameter and a functional on $\mathcal{W}_{0}$, it is not necessarily true that $\mathop{\mathrm{cd}}(G)=\mathop{\mathrm{cd}}(W_{G})$ or $\overline{\mathop{\mathrm{cd}}}(G)=\overline{\mathop{\mathrm{cd}}}(W_{G})$ (see Theorem 4.3). Also, while it is not completely evident from the definitions that $\overline{\mathop{\mathrm{cd}}}(W)\leq\mathop{\mathrm{cd}}(W)$, this inequality does hold for every $W\in\mathcal{W}_{0}$.
It is natural to ask whether for every graphon $W$ there is an $f\in\mathcal{F}$ such that $\overline{\mathop{\mathrm{cr}}}(W)=\overline{\mathop{\mathrm{cr}}}(W,f)$, or a $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$ such that $\mathop{\mathrm{cd}}(W)=\mathop{\mathrm{cd}}(W,\mathcal{D})$ (in other words, can the infimum be substituted by a minimum in the above definitions?). Unfortunately, we do not know the answer to either of these question. The following result, however, is an immediate consequence of the Banach-Alaoglu theorem.
@@
Hence, setting $\overline{\mathcal{I}}$ to be the set formed by all elements of $A$ and all sequential limit points of $A$ with respect to the weak* topology which have image in $[0,1]$ does the job. In the same way, we can show that the set $\mathcal{I}$ which consists of all elements of $B$ and all of its sequential limit points with image in $[0,1]$ has the required properties. ∎
For every $W\in\mathcal{W}_{0}$, let $\overline{I}_{W}\in\overline{\mathcal{I}}$ and $I_{W}\in\mathcal{I}$ denote two functions that attain the minimums in the statement of Lemma 4.1. The questions above can now be restated as follows. Given a graphon $W$, can the functions $\overline{I}_{W}$ and $I_{W}$ always be chosen so that $\overline{I}_{W}=I_{f}$ for some $f\in\mathcal{F}$ and $I_{W}=I_{\mathcal{D}}$ for some $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$? We will come back to this in Section 5.
4.2 Continuity with respect to $d_{\square}$ and connection to graphs
Here, we prove that the crossing and rectilinear crossing densities are continuous with respect to $d_{\square}$, and that they extend the corresponding graph parameters in a precise sense (see Theorem 4.4). We must first establish a weaker sort of continuity for cd and $\overline{\text{cd}}$.
Lemma 4.2.
@@ 4.2 Continuity with respect to $d_{\square}$ and connection to graphs
$\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W)\text{ and }\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W).$
@@
The tensor product of two graphons $W_{1}$ and $W_{2}$ is the function $W_{1}\otimes W_{2}:[0,1]^{4}\rightarrow[0,1]$ defined by $(W_{1}\otimes W_{2})(x,y,z,w)=W_{1}(x,y)W_{2}(z,w)$.
Consider a sequence of graphons $W_{1},W_{2},\dots$ that converges to $W\in\mathcal{W}_{0}$ almost everywhere. Clearly, the sequence $W_{1}\otimes W_{1},W_{2}\otimes W_{2},\dots$ converges to $W\otimes W$ almost everywhere. Choose a subsequence $W_{1}^{\prime},W_{2}^{\prime},\dots$ of $W_{1},W_{2},\dots$ with $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n}^{\prime})=\liminf_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})$ and, for every $n$, let $f_{n}$ be an element of $\mathcal{F}$ such that $\overline{\mathop{\mathrm{cd}}}(W_{n},f_{n})\leq\overline{\mathop{\mathrm{cd}}}(W_{n})+1/2^{n}$. By Banach-Alaoglu, we can pass to a further subsequence $W_{s(1)}^{\prime},W_{s(2)}^{\prime},\dots$ of $W_{1}^{\prime},W_{2}^{\prime},\dots$ such that $\overline{I}_{f_{s(1)}},\overline{I}_{f_{s(2)}},\dots$ converges to some $I\in\overline{\mathcal{I}}$ in the weak* topology. This way,
@@
for every positive integer $n$, and so the expression in the middle converges to
@@
as $n$ goes to infinity.
On the other hand, a standard argument from analysis shows that the sequence $(W_{s(1)}\otimes W_{s(1)})\cdot\overline{I}_{f_{s(1)}},(W_{s(1)}\otimes W_{s(1)})\cdot\overline{I}_{f_{s(1)}},\dots$ converges to $(W\otimes W)\cdot I$ in the weak* topology. Putting everything together, we get that
@@
$\overline{\mathop{\mathrm{cd}}}(W_{n})\leq\frac{1}{8}\int_{[0,1]^{4}}(W_{n}\otimes W_{n})\overline{I}_{W}(x_{1},x_{2},x_{3},x_{4})\prod_{i=1}^{4}dx_{i}$
for every $n$. Since the expression on the right converges to $\overline{\mathop{\mathrm{cd}}}(W)$ as $n$ goes to infinity, we arrive at $\limsup_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\leq\overline{\mathop{\mathrm{cd}}}(W)$, but we also had $\liminf_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\geq\overline{\mathop{\mathrm{cd}}}(W)$; this can only occur if $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W)$.
The fact that $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$ follows verbatim. ∎
Next, we establish a connection between the crossing densities of graphons and the crossing densities of graphs.
@@
For every edge-weighted graph $G$ on $n$ vertices, we have that
$0\leq\overline{\mathop{\mathrm{cd}}}(W_{G})-\overline{\mathop{\mathrm{cd}}}(G)\leq 1/n\text{ and }0\leq\mathop{\mathrm{cd}}(W_{G})-\mathop{\mathrm{cd}}(G)\leq 1/n.$
@@
Let $v_{1},v_{2},\dots,v_{n}$ denote the vertices of $G$ and $I_{1},I_{2},\dots,I_{n}$ be the corresponding subintervals of $[0,1]$ (see the definition of $W_{G}$ in Section 2.2). The argument is similar to the one in the proof of Theorem 3.4. We begin by showing the inequalities for rectilinear crossing densities.
Given any $f\in\mathcal{F}$, we obtain a drawing $\mathcal{D}_{f}$ of $G$ as follows: For every $v_{i}$, choose uniformly at random an element $x_{i}$ from $I_{i}$; the point $f(x_{i})$ will represent $v_{i}$. Draw every edge of $G$ as a segment joining its endpoints and, if necessary, perturb the vertices slightly so that no three of them lie on a line and no three edges have a common interior point. Note that the vertex perturbation step can be carried out without creating new crossings. A straightforward computation shows that
$\mathbb{E}[\overline{\mathop{\mathrm{cd}}}(G,\mathcal{D}_{f})]\leq\overline{\mathop{\mathrm{cd}}}(W_{G},f).$
Since this holds for any $f\in\mathcal{F}$, we get that $\overline{\mathop{\mathrm{cd}}}(G)\leq\overline{\mathop{\mathrm{cd}}}(W_{G})$.
Suppose now that we are given a rectilinear drawing $\mathcal{D}$ of $G$. This drawing induces an $f_{\mathcal{D}}\in\mathcal{F}$, which is defined by setting $f_{\mathcal{D}}(x)=p_{i}$, where $i$ is such that $x\in I_{i}$ and $p_{i}$ is the point that represents $v_{i}$ in $\mathcal{D}$. The value of $\overline{\mathop{\mathrm{cd}}}(f_{\mathcal{D}},W_{G})$ is upper bounded by $\overline{\mathop{\mathrm{cd}}}(G)+1/n$, where the $1/n$ accounts for the crossings coming from the $4$-tuples of points in $[0,1]$ which contain two elements in the same $S_{i}$ (these $4$-tuples are similar to the bad $4$-tuples in the proof of Theorem 3.4). This yields $\overline{\mathop{\mathrm{cd}}}(W_{G})\leq\overline{\mathop{\mathrm{cd}}}(G)+1/n$.
@@
$\mathbb{E}[\mathop{\mathrm{cd}}(G,\mathcal{D}^{\prime})]\leq\mathop{\mathrm{cd}}(W_{G},\mathcal{D}),$
@@
Finally, for every drawing $\mathcal{D}$ of $G$ we obtain an element $\mathcal{D}^{\prime}$ of $\mathcal{D}_{\mathcal{W}}$ in the following way: For every $x\in I_{i}$, let $\mathcal{D}^{\prime}(x)$ be the point that represents $v_{i}$ in $\mathcal{D}$. For every $x,y\in[0,1]$ with $x\in I_{i},y\in I_{j}$ and $i\neq j$, set $\mathcal{D}^{\prime}(x,y)$ to be the curve that represents the edge $(v_{i},v_{j})$. As in the case of rectilinear crossing densities, this graphon drawing satisfies $\mathop{\mathrm{cd}}(W_{G},\mathcal{D}^{\prime})\leq\mathop{\mathrm{cd}}(G)+1/n$. ∎
Theorems 3.5 and 4.3 can now be combined to obtain the main result of this section.
@@
The crossing density and the rectilinear crossing density are continuous with respect to the cut norm $d_{\square}$. Furthermore, if $G_{1},G_{2},\dots$ is a convergent sequence of edge-weighted graphs and the number of vertices of $G_{n}$ goes to infinity with $n$, then $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(G_{i})$ exists and is equal to $\mathop{\mathrm{cd}}(W)$. The same is true with $\overline{\mathop{\mathrm{cd}}}$ in place of $\mathop{\mathrm{cd}}$.